OpenStudy (anonymous):

-6x+3-5=-6x-2 simplify. show your work

OpenStudy (aravindg):

no further possible

OpenStudy (aravindg):

giv kiss

OpenStudy (anonymous):

-6x+3-5=6x-2, -9x-5=6x-2, ANSWER: -4x=4x

OpenStudy (anonymous):

thank u lauren

OpenStudy (anonymous):

Your welcome!!